OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 I got x = -1.8 and x = 2.2 if you don't know about absolute value, please don't help me. I appreciate the effort but if you don't know what your talking about I get really confused.

OpenStudy (anonymous):

@khurramshahzad

OpenStudy (anonymous):

u r wrong

OpenStudy (anonymous):

put |x+1| =-(x+1) u will get wat chck it ?

OpenStudy (anonymous):

I know that, I tried to solve it a few times but I kept getting the same answer. -5x + 1 = 10 -5x = 9 x = -1.8 -5x + 1 = -10 -5x = -11 x = 2.2

OpenStudy (anonymous):

no no ... i did mistake in 1st step my dear bro

OpenStudy (anonymous):

it will be 5x+5=10

OpenStudy (anonymous):

where did the extra 5 come from and where did the 1 go?

OpenStudy (anonymous):

u have to multiply external 5 with all in the | | braces they act like braces

OpenStudy (anonymous):

ohhhhh, ok hold on a sec

OpenStudy (anonymous):

ok now solve it tell me the answers

OpenStudy (anonymous):

x = -1 and x = 3

OpenStudy (anonymous):

no wrong

OpenStudy (anonymous):

look u have simple mistake

OpenStudy (anonymous):

x = -3 and x = 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u knw how it comes ?

OpenStudy (anonymous):

no, because twords the end of my equations at had -15/-5 = 3 and then 5/-5 = -1....

OpenStudy (anonymous):

no u went wrong .. show u how .. wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

−5|x + 1| = 10 first take |x+1|=(x+1) and see -5(x+1)=10 -5x-5=10 -5x=15 x=-3 ok its one answer

OpenStudy (anonymous):

−5|x + 1| = 10 first take |x+1|=-(x+1) and see -5(-)(x+1)=10 +5x+5=10 +5x=10-5 5x=5 x=1 ok its2nd answer

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

are u clear i have just conern with that

OpenStudy (anonymous):

you said there is just 1 answer but there are 2 answers, all my choices were 2 answers. so now im confused.

OpenStudy (anonymous):

ofcource two how one can be possible if u have to use two options for mod | |

OpenStudy (anonymous):

that's not even English bro...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ok potato it my justification clear to u ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

−5|x + 1| = 10 first take |x+1|=(x+1) and see -5(x+1)=10 -5x-5=10 -5x=15 x=-3 ok its one answer . . . . . −5|x + 1| = 10 first take |x+1|=-(x+1) and see -5(-)(x+1)=10 +5x+5=10 +5x=10-5 5x=5 x=1 ok its2nd answer did u get the two answers 1 and -3 how they come ???